# Personalization of Large Language Models: A Survey

### v1

@@ 8.3 Search
Recently, with their growing capabilities in summarization and instruction following, LLMs have been incorporated into search engines (applications such as the new Bing (Microsoft, 2023) and SearchGPT (OpenAI, 2024)), which can provide an engaging conversational process that can help users find information more effectively (Spatharioti et al., 2023; Joko et al., 2024). Incorporating personalization can further tailor results to individual users’ search histories, interests, and contexts, which can lead to more relevant and efficient search experiences (Bennett et al., 2012; Harvey et al., 2013; Cai et al., 2014; Song et al., 2014; Vu et al., 2014; 2015; Zhou et al., 2021). A large number of works (Dou et al., 2007; Sieg et al., 2007; Carman et al., 2010; Teevan et al., 2011; White et al., 2013) focused on how to better personalize search engines before the emergence of LLMs. In the era of LLMs, Zhou et al. (2024b) propose Cognitive Personalized Search (CoPS), a personalized search model that combines LLMs with a cognitive memory mechanism inspired by human cognition. CoPS utilizes sensory, working, and long-term memory components to efficiently process user interactions and improve search personalization without requiring training data. Besides, Jiang et al. (2024) introduce Collaborative STORM, a system that personalizes search experiences by engaging users in multi-turn search sessions and incorporating their interaction history. However, despite the advancements in LLM-augmented search engines, there are still challenges to be addressed. Notably, Sharma et al. (2024) find that users of LLM-powered search exhibit more biased information querying and opinion polarization compared to conventional web search, especially when the LLM reinforces existing views. This phenomenon, known as the “echo chamber” effect (Pariser, 2011; Sharma et al., 2024; Lazovich, 2023; Garimella et al., 2018), emphasizes the challenges in balancing personalization with the need for diverse and objective information retrieval.
@@ 9 Open Problems & Challenges
9.1 Benchmark and Metrics
@@
Effective benchmarks, combined with comprehensive metrics, are crucial for evaluating various aspects of LLMs, including their ability to personalize outputs. However, existing benchmarks for personalization are largely derived from recommendation systems, where the focus is predominantly on final predictions such as ratings, recommended items, or rankings. These benchmarks often overlook the intermediate processes in LLMs’ output generation, which are critical for assessing whether the output is genuinely personalized. LaMP (Salemi et al., 2023) is one of the few benchmarks that specifically targets the evaluation of LLMs in generating personalized outputs. However, LaMP’s scope is limited to text classification and short, single-turn text generation tasks. It lacks the complexity of real-world interactions, which are essential for applications like personalized AI assistants. This gap highlights the need for new benchmarks that can evaluate LLMs’ personalized output generation in more realistic scenarios. Such benchmarks should also integrate personalization perspectives into other key LLM capabilities, including reasoning, planning, instruction following, and long-context understanding, thereby providing a more holistic evaluation. In addition, there is currently no comprehensive quantitative metric to assess the degree of personalization in LLM-generated outputs. Most existing metrics are task-specific and heavily dependent on the downstream task formulations and the quality of gold labels. As a result, they often fail to capture the diverse dimensions of personalization, such as those illustrated in Figure 4. The recent trend of using LLMs as judges in evaluating various aspects of LLM-generated content, due to their versatile nature, presents a promising approach for personalization assessment. Designing an LLM-as-a-Judge framework with personalized criteria rubrics could offer a more nuanced evaluation of the degree of personalization in LLM outputs. However, this approach remains underexplored, and challenges such as instability and potential biases need to be addressed to make it a reliable evaluation method.
@@ 9.3 Stereotype & Bias Issues of Personalization
The personalization of LLMs introduces significant concerns regarding the amplification and perpetuation of stereotypes and biases (Zhang et al., 2023d; Ziems et al., 2024; Gallegos et al., 2024; Li et al., 2023e). When LLMs generate personalized outputs, they rely on data that may inherently contain societal biases related to gender, race, ethnicity, culture, and other sensitive attributes. Personalization can unintentionally reinforce these biases by tailoring content that aligns with the biased data the models are trained on or the ones provided in the prompt, thus exacerbating the problem. For example, recent research (Gupta et al., 2023; Deshpande et al., 2023; Cheng et al., 2023a; Wan et al., 2023) indicates that assigning a specific persona to LLMs, like that of a “disabled person,” can unexpectedly alter their performance across diverse tasks, including those seemingly unrelated to the assigned identity such as general knowledge or mathematical reasoning. Moreover, the feedback loop created by personalized systems can further entrench biases. As LLMs continue to adapt based on user interactions, they might cater to pre-existing preferences and viewpoints, reducing the opportunity for exposure to diverse or corrective perspectives. This can lead to the deepening of echo chambers, where users are repeatedly exposed to biased or stereotypical information without opportunities for counterbalance. Despite growing efforts to mitigate biases in LLMs (Lu et al., 2020; Han et al., 2021; Ghanbarzadeh et al., 2023; Zhang et al., 2023d), there is a limited number of works on how personalization intersects with these biases. He et al. (2024a) introduce Context Steering (CoS), a training-free method for enhancing personalization and mitigating bias in LLMs at inference time by quantifying and modulating the influence of contextual information on model outputs. CoS works by computing the difference in token prediction likelihoods between models with and without context, then using this difference to adjust the influence of context during generation. Vijjini et al. (2024) introduce the concept of personalization bias in LLMs, where LLM performance varies based on the user’s demographic identity provided during the interaction. It proposes a framework to evaluate and quantify this bias by measuring safety-utility trade-offs across different user identities, using widely-used datasets. The study demonstrates the prevalence of personalization bias in both open-source and closed-source LLMs, and while it explores mitigation strategies like preference tuning and prompt-based defenses, it concludes that completely eliminating this bias remains an open challenge, highlighting the need for further research in this area. Overall, it is critical to design personalization systems that actively account for fairness and inclusivity, ensuring that personalized outputs do not reinforce harmful stereotypes or perpetuate biased perspectives. Future work should explore techniques such as bias detection during the personalization process, incorporating fairness constraints into the personalization pipeline, and ensuring that diverse perspectives are represented in user-specific outputs. Addressing these challenges will require careful consideration of the trade-offs between personalization and fairness, as well as the development of robust evaluation frameworks that measure not only the degree of personalization but also the impact on bias and stereotype propagation. Ultimately, ensuring that personalized LLMs promote inclusivity and fairness is essential to building systems that are socially responsible and beneficial to all users.
@@ 10 Conclusion
This survey provides a unified and comprehensive view of the burgeoning field of personalized LLMs. It bridges the gap between the two dominant lines of research — direct personalized text generation and leveraging personalized LLMs for downstream applications — by proposing a novel taxonomy for personalized LLM usage and formalizing their theoretical foundations. An in-depth analysis of the personalization granularity highlights trade-offs among user-level, persona-level, and global preference alignment approaches, laying the groundwork for future hybrid systems that can dynamically adapt to user needs and data availability. Additionally, this survey provides a detailed examination of techniques for personalizing LLMs, shedding light on the strengths and limitations of each approach. We explore the nuances of retrieval-augmented generation, various prompting strategies, supervised fine-tuning approaches, embedding learning methods, and the evolving landscape of learning from personalized feedback through RLHF, underscoring the need for more robust and nuanced methods. Finally, our comprehensive survey of evaluation methodologies and datasets highlights the critical need for new benchmarks and metrics specifically designed for assessing personalized LLM outputs. Despite the significant progress made in personalized LLMs, numerous challenges and open problems remain. Key areas for future research include addressing the cold-start problem in low-resource scenarios, mitigating stereotypes and biases in personalized outputs, ensuring user privacy throughout the personalization pipeline, and extending personalization to multi-modal systems. The field of personalized LLMs is rapidly evolving, with the potential to revolutionize human-AI interaction across diverse domains. By understanding the foundations, techniques, and challenges outlined in this survey, researchers and practitioners can contribute to the development of more effective, fair, and socially responsible personalized LLMs that cater to diverse user needs and preferences.

### v3

@@ 8.3 Search
Recently, with their growing capabilities in summarization and instruction following, LLMs have been incorporated into search engines (applications such as the new Bing (Microsoft, 2023) and SearchGPT (OpenAI, 2024)), which can provide an engaging conversational process that can help users find information more effectively (Spatharioti et al., 2023; Joko et al., 2024). Incorporating personalization can further tailor results to individual users’ search histories, interests, and contexts, which can lead to more relevant and efficient search experiences (Bennett et al., 2012; Harvey et al., 2013; Cai et al., 2014; Song et al., 2014; Vu et al., 2014; 2015; Zhou et al., 2021). A large number of works (Dou et al., 2007; Sieg et al., 2007; Carman et al., 2010; Teevan et al., 2011; White et al., 2013) focused on how to better personalize search engines before the emergence of LLMs. In the era of LLMs, Zhou et al. (2024b) propose Cognitive Personalized Search (CoPS), a personalized search model that combines LLMs with a cognitive memory mechanism inspired by human cognition. CoPS utilizes sensory, working, and long-term memory components to efficiently process user interactions and improve search personalization without requiring training data. Besides, Jiang et al. (2024b) introduce Collaborative STORM, a system that personalizes search experiences by engaging users in multi-turn search sessions and incorporating their interaction history. However, despite the advancements in LLM-augmented search engines, there are still challenges to be addressed. Notably, Sharma et al. (2024) find that users of LLM-powered search exhibit more biased information querying and opinion polarization compared to conventional web search, especially when the LLM reinforces existing views. This phenomenon, known as the “echo chamber” effect (Pariser, 2011; Sharma et al., 2024; Lazovich, 2023; Garimella et al., 2018), emphasizes the challenges in balancing personalization with the need for diverse and objective information retrieval.
@@ 9 Open Problems & Challenges
9.1 Comprehensive Evaluation: Benchmarks, Automatic Metrics, LLM-as-a-Judge, and Beyond
@@
Effective benchmarks, combined with comprehensive metrics, are crucial for evaluating various aspects of LLMs, including their ability to personalize outputs. However, existing benchmarks for personalization are largely derived from recommendation systems, where the focus is predominantly on final predictions such as ratings, recommended items, or rankings. These benchmarks often overlook the intermediate processes in LLMs’ output generation, which are critical for assessing whether the output is genuinely personalized. LaMP (Salemi et al., 2023) is one of the few benchmarks that specifically targets the evaluation of LLMs in generating personalized outputs. However, LaMP’s scope is limited to text classification and short, single-turn text generation tasks. It lacks the complexity of real-world interactions, which are essential for applications like personalized AI assistants. This gap highlights the need for new benchmarks that can evaluate LLMs’ personalized output generation in more realistic scenarios. Such benchmarks should also integrate personalization perspectives into other key LLM capabilities, including reasoning, planning, instruction following, and long-context understanding, thereby providing a more holistic evaluation. Overall, we envision comprehensive personalization benchmarks that effectively capture multi-turn interactions, evolving user preferences, and diverse contextual scenarios. For future directions in designing personalization benchmarks, one promising area involves language agents, which are increasingly applied to tasks such as web navigation (Deng et al., 2024), scientific discovery (Chen et al., 2024e; Si et al., 2024), research support (Asai et al., 2024), and coding assistance (Wang et al., 2024e). While their evaluation typically focuses on task success rates, the role of personalization in these contexts remains underexplored and calls for the development of dedicated evaluation criteria. Additionally, while current benchmarks predominantly emphasize English, there is an urgent need to expand their scope to include multilingual settings to ensure broader applicability and inclusivity. Given that LLMs are known to be sensitive to prompt variations (Zhuo et al., 2024), addressing distribution shifts and robustness in prompt design, as well as out-of-distribution scenarios (Wang et al., 2023a), becomes essential. Controlled experiments that systematically isolate and analyze dimensions such as style, topic, and user preferences can provide valuable insights. Furthermore, incorporating cultural and values-based adaptations (Shi et al., 2024b), alongside dialectal variations (Ziems et al., 2023), will enable these benchmarks to better reflect the real-world complexity of personalization and its diverse user base. Additionally, evaluating on static benchmarks often entails limitations, such as susceptibility to data contamination (Sainz et al., 2023; Balloccu et al., 2024) and a lack of adaptability to evolving LLM capabilities. To address these issues, designing dynamic evaluation frameworks for personalization tasks offers a promising alternative, enabling assessments with controllable complexity (Zhu et al., 2023b; Zhang et al., 2024c).
In addition, there is currently no comprehensive quantitative metric to assess the degree of personalization in LLM-generated outputs. Most existing metrics are task-specific and heavily dependent on the downstream task formulations and the quality of gold labels. As a result, they often fail to capture the diverse dimensions of personalization, such as those illustrated in Figure 4. The recent trend of using LLMs as judges in evaluating various aspects of LLM-generated content, due to their versatile nature, presents a promising approach for personalization assessment. Designing an LLM-as-a-Judge framework with personalized criteria rubrics could offer a more nuanced evaluation of the degree of personalization in LLM outputs. However, this approach remains underexplored, and challenges such as instability and potential biases need to be addressed to make it a reliable evaluation method. Specifically, challenges arise when LLM-based evaluators unintentionally inject their own internal persona or biases (Jiang et al., 2024a), and they can sometimes exhibit overconfidence by offering overly favorable assessments of their own outputs (Tian et al., 2023). Consistency is another issue (Li et al., 2024c), as LLM-as-a-judge evaluations often fluctuate depending on the context (Gu et al., 2024) and exhibit phenomena such as positional biases (Li et al., 2024i), which poses significant problems in personalization tasks involving nuanced user attributes. There are also concerns about robustness, given that adversarial inputs can undermine the integrity of LLM-based evaluations (Doddapaneni et al., 2024; Raina et al., 2024; Shi et al., 2024a; Zheng et al., 2024). Furthermore, efficiency and flexibility can be limited when LLM-based evaluators must be manually prompted for each new personalization scenario. Despite these challenges, there are significant opportunities for improvement. Developing more refined prompting formats and strategies, such as scalar scoring, pairwise comparisons, and multiple-choice selections, can provide more nuanced and reliable evaluations. Additionally, addressing open problems like in-context exemplar selection could further enhance the evaluation process. Given the computational overhead of large-scale LLM-as-a-judge frameworks, which often rely on extensive API calls for model inference (Gu et al., 2024), training specialized smaller models tailored for personalization evaluation presents a promising alternative for reducing costs and improving efficiency (Kim et al., 2024; Huang et al., 2024a). Furthermore, exploring agent-as-a-judge paradigms (Zhuge et al., 2024) that incorporate multi-agent collaboration, tool integration, or human-in-the-loop approaches offers a path toward greater transparency, fairness, and robustness in assessing personalized outputs.
In summary, evaluating robust personalization in LLMs presents unique challenges compared to the evaluation of other capabilities. Personalization objectives in real-life applications are inherently diverse, leading to pluralistic alignment requirements (Sorensen et al., 2024). The varying levels of personalization defy reliance on any single evaluation criterion. Furthermore, the inherent ambiguity and subjectivity of user preferences add complexity, underscoring the urgent need for a finer-grained taxonomy to capture the full spectrum of personalization phenomena. Current datasets often suffer from imbalanced data, which can lead to incomplete or biased assessments of personalization performance. For instance, datasets may overrepresent users with frequent online activity, potentially skewing models to prioritize the preferences and behaviors of these users while underrepresenting less active user groups. Additionally, user preferences are dynamic and can shift over time, necessitating the development of multi-turn benchmarks that reflect evolving personalization goals and longitudinal variations. Moreover, due to inconsistencies in dataset structures, task formulations, and the diverse scenarios in real-life applications, direct performance comparisons can be challenging, making it difficult to achieve a unified and fair evaluation of different techniques. It is also critical to address privacy, bias, and fairness—factors that may conflict with certain personalization objectives—to ensure personalized systems remain safe, ethical, and equitable. Given these challenges, we advocate for collaborative efforts to create open-source platforms and shared tasks that allow personalized LLMs to be tested against standardized metrics aligned with core LLM capabilities. Finally, we emphasize the importance of cross-domain generalization tests beyond recommendation systems and text generation to evaluate how well personalization systems adapt across diverse domains and contexts.
@@ 9.3 Stereotype & Bias Issues of Personalization
The personalization of LLMs introduces significant concerns regarding the amplification and perpetuation of stereotypes and biases (Zhang et al., 2023d; Ziems et al., 2024; Gallegos et al., 2024; Li et al., 2023f). When LLMs generate personalized outputs, they rely on data that may inherently contain societal biases related to gender, race, ethnicity, culture, and other sensitive attributes. Personalization can unintentionally reinforce these biases by tailoring content that aligns with the biased data the models are trained on or the ones provided in the prompt, thus exacerbating the problem. For example, recent research (Gupta et al., 2023; Deshpande et al., 2023; Cheng et al., 2023a; Wan et al., 2023) indicates that assigning a specific persona to LLMs, like that of a “disabled person,” can unexpectedly alter their performance across diverse tasks, including those seemingly unrelated to the assigned identity such as general knowledge or mathematical reasoning. Moreover, the feedback loop created by personalized systems can further entrench biases. Besides, Weissburg et al. (2024) examines bias in LLMs as personalized educators, showing disparities in content selection and difficulty based on demographics like race, gender, income, and disability. Using 17,000+ explanations and two bias metrics, it finds all tested LLMs exhibit bias, with income and disability status most affected. As LLMs continue to adapt based on user interactions, they might cater to pre-existing preferences and viewpoints, reducing the opportunity for exposure to diverse or corrective perspectives. This can lead to the deepening of echo chambers, where users are repeatedly exposed to biased or stereotypical information without opportunities for counterbalance. For example, Kantharuban et al. (2024) investigates how large language models generate recommendations that reflect both explicit and implicit cues of a user’s identity, leading to biased outputs that align with racial stereotypes. The authors show that although personalization can tailor recommendations to the user’s background, it often restricts the range of options for minority users and obscures the role of identity in the recommendation process. Despite growing efforts to mitigate biases in LLMs (Lu et al., 2020; Han et al., 2021; Ghanbarzadeh et al., 2023; Zhang et al., 2023d), there is a limited number of works on how personalization intersects with these biases. He et al. (2024a) introduce Context Steering (CoS), a training-free method for enhancing personalization and mitigating bias in LLMs at inference time by quantifying and modulating the influence of contextual information on model outputs. CoS works by computing the difference in token prediction likelihoods between models with and without context, then using this difference to adjust the influence of context during generation. However, CoS does not specifically aim to mitigate biases that arise within the personalization process. Vijjini et al. (2024) introduce the concept of personalization bias in LLMs, where LLM performance varies based on the user’s demographic identity provided during the interaction. It proposes a framework to evaluate and quantify this bias by measuring safety-utility trade-offs across different user identities, using widely-used datasets. The study demonstrates the prevalence of personalization bias in both open-source and closed-source LLMs, and while it explores mitigation strategies like preference tuning and prompt-based defenses, it concludes that completely eliminating this bias remains an open challenge, highlighting the need for further research in this area. Overall, it is critical to design personalization systems that actively account for fairness and inclusivity, ensuring that personalized outputs do not reinforce harmful stereotypes or perpetuate biased perspectives. Future work should explore techniques such as bias detection during the personalization process, incorporating fairness constraints into the personalization pipeline, and ensuring that diverse perspectives are represented in user-specific outputs. For example, general LLM debiasing methods can serve as inspiration, including data-centric approaches such as data augmentation, filtering, reweighting training data, or modifying loss functions during fine-tuning to directly mitigate biases. Additionally, inference-time techniques—such as adjusting decoding strategies or rewriting model outputs—can help reduce bias in personalized responses while maintaining the model’s tailored behavior (Gallegos et al., 2024). Addressing these challenges will require careful consideration of the trade-offs between personalization and fairness, as well as the development of robust evaluation frameworks that measure not only the degree of personalization but also the impact on bias and stereotype propagation. Ultimately, ensuring that personalized LLMs promote inclusivity and fairness is essential to building systems that are socially responsible and beneficial to all users.
@@ 9.5 Multimodality
Specifically, according to Wu et al. (2024a), the unique high-level challenges in personalized multimodal LLM systems can be categorized as follows. (1) The integration of heterogeneous data. LMMs must process and align diverse data types such as text, images, audio, and structured user interaction history (Li et al., 2024f). Encoding discrepancies arise as different modalities require distinct encoding techniques, making fusion difficult. Additionally, cross-modal alignment remains an issue since text and images may contain complementary but sometimes conflicting information. Another difficulty is data sparsity, where some users interact predominantly with one modality, leading to imbalanced personalization. Future opportunities include developing unified embedding spaces to bridge the gap between diverse modalities, leveraging self-supervised learning to enhance cross-modal understanding, and designing adaptive models that dynamically adjust weights for each modality based on user interaction patterns. (2) Data noise, redundancy, and quality control. Different modalities often include noisy, redundant, or irrelevant information (Liu et al., 2024e; Lyu & Luo, 2022). For example, images of the same object may vary in quality, while textual descriptions may be verbose or contain unnecessary details. Extracting meaningful insights while filtering out redundant or noisy data is essential for effective personalization. Future work should focus on implementing noise-aware training strategies to filter out irrelevant data during model training, using multimodal attention mechanisms to prioritize relevant user interactions and suppress redundant information, and employing contrastive learning to improve robustness against low-quality or inconsistent data. (3) Granular understanding of multimodal data. While text-based LLMs excel at linguistic processing, capturing subtle visual or auditory cues remains difficult (Shen et al., 2024). User preferences in areas such as fashion, art, or music often depend on fine-grained multimodal details, such as color, texture, or rhythm. Personalized multimodal models must enhance their ability to extract and relate these details meaningfully across different input types. Future advancements should develop hierarchical representations that preserve both fine-grained and high-level information, improve multimodal contrastive learning to align visual, auditory, and textual representations effectively, and explore fine-tuned retrieval-based techniques to improve nuanced personalization.
Beyond these overarching challenges, specific tasks introduce unique obstacles. In personalized image generation, achieving a balance between fidelity and diversity remains a key challenge. Hybrid diffusion models have shown promise in addressing this issue by refining subject representation and controlling generation constraints (Wang et al., 2024c; Ma et al., 2024; Song et al., 2025). Additionally, tokenization presents another challenge, as indexing multimodal inputs into lookup tables for efficient personalized generation requires specialized designs (Gal et al., 2022). For multi-modal personalized recommendation, multi-modal collaborative filtering introduces additional complexities. Recent approaches have attempted to unify multi-channel information, integrating generative recommendation mechanisms alongside dynamic item modifications (Yu et al., 2024). These methods enable tokenization of both items and user embeddings, facilitating multimodal feature incorporation into the model’s latent space. Addressing these challenges will require continued research into improved multimodal alignment, data filtering, and scalable architectures that enhance personalization across diverse applications. Future advancements in integrating multimodal feedback loops and adaptive learning mechanisms will be essential for unlocking the full potential of personalized multimodal LLMs.
@@ 10 Conclusion
This survey provides a unified and comprehensive view of the burgeoning field of personalized LLMs. It bridges the gap between the two dominant lines of research — direct personalized text generation and leveraging personalized LLMs for downstream applications — by proposing a novel taxonomy for personalized LLM usage and formalizing their theoretical foundations. An in-depth analysis of the personalization granularity highlights trade-offs among user-level, persona-level, and global preference alignment approaches, laying the groundwork for future hybrid systems that can dynamically adapt to user needs and data availability. Additionally, this survey provides a detailed examination of techniques for personalizing LLMs, shedding light on the strengths and limitations of each approach. We explore the nuances of retrieval-augmented generation, various prompting strategies, representation learning approaches, and the evolving landscape of learning from personalized feedback through RLHF, underscoring the need for more robust and nuanced methods. Finally, our comprehensive survey of evaluation methodologies and datasets highlights the critical need for new benchmarks and metrics specifically designed for assessing personalized LLM outputs. Despite the significant progress made in personalized LLMs, numerous challenges and open problems remain. Key areas for future research include addressing the cold-start problem in low-resource scenarios, mitigating stereotypes and biases in personalized outputs, ensuring user privacy throughout the personalization pipeline, and extending personalization to multi-modal systems. The field of personalized LLMs is rapidly evolving, with the potential to revolutionize human-AI interaction across diverse domains. By understanding the foundations, techniques, and challenges outlined in this survey, researchers and practitioners can contribute to the development of more effective, fair, and socially responsible personalized LLMs that cater to diverse user needs and preferences.